The distance. S metres, in which a car can stop is related to its speed, 'km/h by
S=pV+ qv² where p and q are constants. A car travelling at 10 km/h can stop in a distance of 5 metres and at 20 km/h in a distance of 12 metres. Calculate the distance in which a car travelling at 30 km/h can stop

Answer by ikleyn(52780)

.
he distance. S metres, in which a car can stop is related to its speed, 'km/h by
S=pV+ qv² where p and q are constants. A car travelling at 10 km/h can stop in a distance of 5 metres
and at 20 km/h in a distance of 12 metres. Calculate the distance in which a car travelling at 30 km/h can stop
~~~~~~~~~~~~~~~~~~~~~~~~~~

As you read the problem, write two equations

    10*p + 10^2*q =  5  meters  (for the speed V = 10 km/h)

    20*p + 20^2*q = 12  meters  (for the speed V = 20 km/h)


Or, the same

    10p + 100q =  5    (1)

    20p + 400q = 12    (2)


To solve these equations, multiply equation (1) by 2 (both sides).  Keep equation (2) as is

    20p + 200q = 10    (1')

    20p + 400q = 12    (2')


From equation (2'), subtract equation (1').  You will get

          200q =  2

             q = 2/200 = 0.01.


Then from equation (1)

    10p + 100*0.01 = 5

    10p = 5 - 1 = 4

      p = 4/10 = 0.4.


Thus, now you know p= 0.4, q= 0.01.


Hence, at V= 30 km/h,  the distance to stop is

   30p + 30^2*q = 30*0.4 + 900*0.01 = 21,


ANSWER.  The distance to stop at V 30 km/h is 21 meter.

Solved.